Can these systems be solved with the addition method. I am not getting it if it can. Show steps so I may see where I went wrong. Please thank you.
3x - 2y =8
x = -2y
and 4y = -5x
5x + 8y = 20
Thank you

Answer by stanbon(75887)

3x - 2y =8
x = -2y
---
Rearrange:
3x - 2y = 8
x + 2y = 0
--------------
Add to solve for "x":
4x = 8
x = 2
----------
Subsitute to solve for "Y":
2 = -2y
y = -1
==============
4y = -5x
5x + 8y = 20
---
Rearrange:
-5x - 4y = 0
5x + 8y= 20
----------------
Add to get:
4y = 20
y = 5
-----
Substitute to solve for "x":
-5x = 4*5
x = -4
=======================
Cheers,
Stan H.